OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3^x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

550=35*3^x 550/35=3^x \[\log_3(\frac{550}{35})=x\]

OpenStudy (anonymous):

about 3.5 hours? 2.5, 3.5?

OpenStudy (anonymous):

opps i mean 5.5